OpenStudy (anonymous):

Nautical speed was once measured by throwing a rope into the water from the ship. The rope had knots every 47 ft 3 in, and its wedge-shaped end would "grab" the water. The speed was the number of rope knots that went into the water while a 28-second hourglass ran down. A nautical mile is about 6076 feet. What is the ratio of a nautical mile to the length of the rope between knots?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

could you check if it is either 128.6 times as long 129.3 times as long 129.8 times as long 128.1 times as long

OpenStudy (anonymous):

well a knot is 47 ft 3 in this can be rewritten fractionally as 47 ft + 1/4 ft = 47.25 ft (a foot is 12 inches so 3 inches is 3/12 ft or 1/4 ft) now you want the ratio between a nautical mile and a knot all you have to do here is divide! 6076/47.25 = ? to simplify your fraction, you can multiply the top and bottom by 4 6076*4 = 24304 47.25*4 = 189 so you want to simplify 24304/189... this'll get pretty ugly pretty fast.. but unless otherwise specified, I'd leave it as this!

OpenStudy (anonymous):

thanks but the answers are either 128.6 times as long 129.3 times as long 129.8 times as long 128.1 times as long

OpenStudy (anonymous):

i tried all of them and they dont work

OpenStudy (anonymous):

i'll double check my math really quick...

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

It should be 128.6 (128.59 rounds up) But the numbers I gave you worked! Either use a calculator, or use long division for the first 5 numbers!

OpenStudy (anonymous):

gracias i will try to the problem again so i can practice